OpenStudy (anonymous):

Use algebra to find each point at which the line (x-2y=-5) intersects the circle (x^2 + y^2 = 10)

OpenStudy (anonymous):

taking line x-2y=-5 x = 2y-5 substitute this x in circle equation then u will get a quadratic equation, simplify it

OpenStudy (anonymous):

you will get 2 x values again by substituting x values in line equation get corresponding y values you will now have two pairs of points (x1,y1) and (x2,y2) thats it!

OpenStudy (anonymous):

x= 2y+5 circle x^ +y^=10 2y+5^ +y^=10 4y^+20y+25+y^=10 where does 20y come from sorry 10 years since i did math

OpenStudy (anonymous):

(2y+5)(2y+5)=2y*2y+2y*5+5*2y+5*5=4y^2+10y+10y+25=4y^2+20y+25